Project

General

Profile

News

Executive Board: Meeting 9 July 2011 (1 comment)

Added by Tunnell, Christopher almost 11 years ago

Agenda:

  1. Actions + minutes
  2. Schedule slide FAC
  3. Software
  4. Publications
  5. Conference

MICE: MICE News

Added by Kyberd, Paul about 11 years ago

Report from the Target Group

During running on Wednesday 13 July, the target in ISIS, which had previously
been operating normally, performed two actuations where it was observed
to hit the lower stop rather than returning to "hold" position. Shortly
afterwards, it overshot the hold position and went too high. This
behaviour has not been observed before, and is thought to be due to a
malfunction of the position-measuring system.

The target frame has now been raised so that the target cannot intercept
the beam. Further investigations are in progress, but it appears likely
that the target has jammed in the upper bearing and cannot be
actuated. If this is the case, the target will be out of action until
it can be replaced in the next ISIS shut-down".

Implications of the target failure: Spokesman

This target has successfully performed 571k pulses and it had been
decided in agreement with ISIS to leave it in the ring until it broke
while better design was being developed. Both of these things have now
happened. We are fortunate enough to have at least one design that works
(and an even better one in preparation).

This means that the summer run had to be canceled, something that caused
disappointment -- we all like to see MICE alive and running!
Nevertheless the EMR test, which was the main aim of the run, could be
sufficiently successful in the few days of running starting 29 June. The
EMR will therefore be sent back to University of Geneva.

The baseline schedule is that we re-install a new target before the
planned running in February-March 2012 (full EMR run). If anyone has a
strongly motivated proposal for data taking before then please let the
MOM and Spokesperson know as soon as possible.

There is no reason to believe that this incident will have an impact on
the schedule that was agreed at the collaboration meeting (EMR run in
February 2012, StepIV run in October 2012 to March 2013).

August Bank Holiday at RAL: Friday 26th – Monday 29th August:

· R22 closes at 4pm on Fri 26th Aug and re-opens at 07:30 on Tuesday 30th August
· R1 Coffee Lounge closes at 4pm on Fri 26th Aug and re-opens at 08:30 on Wednesday 31st August
· Electron cafe closes at 3:30pm on Fri 26th Aug and re-opens at 08:30 on Tuesday 30th August
· Breakfast will be available in Ridgeway House for anyone who is at RAL over the weekend

Reduced service on Tuesday 30 August 2011:
· R22 open for breakfast and lunch only, from 07:30 to 16:00 (there will be no dinner service)
· Electron Cafe open, from 08:30 to 15:30
· R1 Coffee Lounge closed

Both beams are down and any guests in staying in Ridgeway will be advised on alternative venues for dinner.

Executive Board: Meeting 17 June 2011 (1 comment)

Added by Tunnell, Christopher about 11 years ago

Main items:
  • Minutes and actions
  • Schedule discussion (Andy Nichols)
  • MOU (Maurizio et al)
  • MICE step I papers status (Maurizio)
  • Summer conferences
  • a o b

MICE: MICE news 19th June

Added by Kyberd, Paul about 11 years ago

MICE News - 19th June 2011

TOF1 has been re-installed at RAL after refurbishing in Milano Bicocca
in the week 13/6-17/6. Unfortunately not all PMTs have been exchanged, as
some to be refurbished are still in Japan at Hamamatsu, due to delays
from the last Tsunami. The EMR support nicely fitted inside the
Rm3 downstream platform. So both TOF2, KL and the first planes of EMR
are in place for the next data taking.

MICE: MICE news May 30 2011

Added by Kyberd, Paul about 11 years ago

Mice Special news from the Spokesmouse

MICE Schedule Review

As I announced in the MICE video conference, in preparation for the MICE project board of 28 June, we held 23-25 May an internal schedule review of MICE.

The agenda with presentations can be found here https://indico.cern.ch/conferenceDisplay.py?confId=136454 Notes and the recommendations by external reviewers will be posted asap.

Please note that this is an internal review and that the information contained in the talks is not "MICE-official"!

The review was very useful, was quite interactive, raised a number of important questions and provided many elements. The information thus gathered will allow to re-baseline the schedule, at least to step IV. Here are some (selected) highlights. <p>

Step I
1. The EMR test run in July is confirmed, it has been proposed to delay the full EMR data
taking to the winter period (February-March 2012)

Step IV
1. There is now a schedule and milestones for the spectrometer solenoids, the manpower has been considerably reinforced, we are only waiting for the final decision on the quench protection system to start the clock ticking.
2. The first AFC magnet has been wound and a schedule exists for commissioning at RAL in the fall of this year (2011).
3. the LH2 absorber number one is on its way to RAL from Japan.
4. the LH2 R&D is progressing well in the MICE hall, next milestone in July 2011 with LHe filling, final results by the end of 2011.
5. the solid absorbers are getting defined and the diffuser is in construction at Oxford.
6. and as can be expected there is some head-scratching to do to dispatch efficiently all
these activities between the MICE hall and the R9 building at RAL.
7. Controls and monitoring are taking shape -- this is becoming a real cornerstone and urgency of the experiment, volunteers are more than welcome to join, please get in touch with Pierrick Hanlet.<br>
8. Software organization is evolving to a more structured and controlled environment --
help from every involved MICE is important.

Taking all into account, this will lead to a new schedule of MICE step IV by mid-june, in time for the MPB on 28 June and final approval by MICE at the next collaboration meeting.

STEP V-VI
1. The Coupling coils are the next priority of the LBNL group, who are presently actively
searching a large cryostat to test the first, already wound, full size coil.
2. There was considerable progress in the design of the layout of RF power in the hall.
3. We should work hard to be in position to have achieved data taking to demonstrate ionization cooling before the next long ISIS shut down (8 months) starting in August 2014.

Again, many thanks to all contributors, and in particular to Andy Nichols who masterminded the operation.

New MICE online Group Coordinator

Linda Coney as new MICE online group (MOG) coordinator

It is a pleasure to inform you all MICE that Linda has agreed to take
on the co-ordination of the online group in replacement of
Jean-Sebastien who is leaving MICE for another life in July. Let us take
this opportunity to thank "JSG" for having opened the trail and led the
way for the last 3 years, and the online group for ensuring the success
of data taking for step I. The Online group will remain a cornerstone of
MICE activities until the last muon has made it though MICE step VI.

Everyone is enthusiastic about Linda and we are convinced that all will
push in the same direction to help the group achieve its missions!

MICE: MICE News. May 22

Added by Kyberd, Paul about 11 years ago

TOF2-KL
The changes in the TOF2-KL trolley have been made as designed
and it's ready for the EMR integration. <br>
TOF2 and KL detectors have been reinstalled along the beam line
and all PMTs checked.

MICE: MICE News 15 May

Added by Kyberd, Paul about 11 years ago

News from Tesla
The winding of the two coils on the bobbin for the first Absorber and Focus Coil Cryostat has been completed at Tesla Engineering.

Cerenkov
D. Sanders and L. Cremaldi will visit RAL in mid-June to work on CKOV detector and interact with the MICE software groups.

Power Status in the MICE Local Control Room
DAQ machines OFF
PPS rack OFF
Target rack OFF
DS rack OFF
Magnet rack OFF
RAID machines OFF
ONREC01 OFF
OPIPC1 & 3 OFF
MICESTORE OFF

Network switches ON UPS on wall power
IOCPC1 ON wall
ONREC02 ON wall
ECSERV2 ON wall
IOCPC2 ON wall
ECSERV ON wall
CAGATEWAY ON wall
OPIPC2 ON wall

Checked rack room and all consumer units isolated OFF
Checked all hall racks and all consumer units isolated OFF

Executive Board: Meeting 20 May 2011 (2 comments)

Added by Tunnell, Christopher about 11 years ago

Agenda:

  • minutes of last meeting and review of actions
  • matters arising from TB and solenoid workshop (Nichols)
  • MICE schedule review status (Andy Nichols)
  • MICE running and MOM modifications :Roumen Tsenov asks to be replaced and proposes Galina see below
  • matters arising from VC (all)
  • MICE Paper I report (see below)
  • Editorial board (Maurizio)

*update of MICE note 95 including analysis of data (Alain will circulate proposal independently)

  • agenda of meeting in July (Oxford) (John Cobb)
  • Common fund (Ken Long)

*In absence of Mike Zisman we can have a brief discussion on the common fund proposal

  • Speakers board (Vittorio)
  • A.o.B.

Executive Board: Meeting 15 April 2011 (1 comment)

Added by Tunnell, Christopher about 11 years ago

Agenda:
  • Minutes of last meeting and review of actions
  • Matters arising from the TB (Nichols)
  • MICE schedule review status (Nichols)
  • MICE running
  • Matters arising from the VC (all)
  • Common Fund and request from USA (Zisman, Long)
  • MICE Paper I (Cobb)
  • Agenda of meeting in July (Oxford) (Cobb)
  • Common fund (Long)
  • Editorial board (Bonesini)
  • Speakers board (Palladino)
  • AOB
(111-120/140)

Also available in: Atom